# DND: Boosting Large Language Models with Dynamic Nested Depth

### v1

@@ Abstract
We introduce Dynamic Nested Depth (DND), a novel method that improves performance for off-the-shelf LLMs by selecting critical tokens to reprocess in a nested depth manner. Specifically, at the end of the given transformer layer, DND identifies more critical tokens with a router and feeds them back for an extra round of processing, effectively “reviewing” difficult tokens while avoiding redundant computation for easier ones. The dynamic selection mechanism is tailored for precise control via two novel strategies: a router controlling loss to enhance token selection distinguishability, and a threshold control scheme to ensure selection stability. We demonstrate the effectiveness of DND by directly integrating it into pre-trained dense and MoE models during a post-training phase. On diverse benchmarks, this approach boosts the performances of the dense Qwen3-1.7B by 1.88% and the MoE Qwen3-30B-A3B by 0.87%, all with a minimal parameter and computing increase.
@@
Large Language Models (LLMs) have transformed artificial intelligence with their powerful abilities. The main strategy for improving them has been scaling, as empirical laws show that model performance predictably increases with more parameters, data, and computation (Achiam et al., 2023, Team et al., 2024, Yang et al., 2025, Liu et al., 2024). However, this scaling paradigm has significant drawbacks. The computational overhead for both training and inference grows exponentially with model size. This trend underscores a critical need for more efficient approaches to enhance model performance beyond simple brute-force scaling.
A key insight from (Gloeckle et al., 2024) is that prediction difficulty varies significantly across tokens; some are trivial to predict, while others demand deep computational processing. This disparity motivates token-level adaptive computation, where models can focus resources on the most critical inputs. A foundational version of this approach is token pruning, which has been shown to be effective across language understanding (Bae et al., 2025), model compression (Yang et al., 2022b), and vision (Hojjat et al., 2025). By filtering out uninformative tokens, these methods reduce computational overhead and can even improve robustness by mitigating noise. This establishes a binary choice: a token is either discarded or processed normally. Furthermore, we propose the natural next step: instead of merely retaining challenging tokens for standard processing, we should allocate additional computation to them, ensuring these critical tokens are properly understood.
Our choice of how to allocate additional computation is inspired by latent strategies in test-time scaling (Hao et al., 2024, Saunshi et al., 2025). Unlike using explicit output expansion like COT, these methods recur computation in hidden states, scaling inference without extra text generation. They observed that reasoning tokens place uneven demands on computation: as illustrated in Fig. 1, most tokens serve fluency, while a few critical ones drive complex planning or logical transitions.
@@
To achieve dynamic selection and recalculation of critical tokens, we propose a novel architecture and training strategy. Specifically, as shown on the right of Fig. 1, we incorporate a linear layer as a router at the end of the transformer block with DND strategies. To achieve more robust routing and avoid potential information leakage during the inference of auto-regressive LLMs (Raposo et al., 2024), we adopt a token-choice routing strategy. In this approach, each token is routed independently based on whether its output exceeds a predefined threshold. The selected tokens are reorganized into a new sequence and are re-fed to compute the dynamic nested output. Moreover, as DND is a post-training method, we carefully design a normalized fusion strategy that integrates the dynamic nested output with the original forward output to preserve global pre-training knowledge.
In terms of training strategies, since our routing method treats each token independently, it lacks the precision of the top-k ratio routing when selecting tokens for recurrent computation (Raposo et al., 2024). To address this, we carefully design training strategies to control both the routers that determine token selection preferences and the thresholds that ultimately decide whether a token is selected. To enhance token selection distinguishability, we optimize the router’s output distribution, encouraging the outputs across tokens to be distinguishable through a router controlling loss. To stabilize the token selection ratio during training, we adopt a threshold control scheme, where the threshold is updated based on the error between the expected ratio and the actual ratio computed over a sample buffer. Furthermore, to ensure more synchronized control, we update the threshold by blending it with the average top-k routing values using an exponential moving average (EMA).
Our method is a post-training approach that can be directly integrated into existing dense and Mixture-of-Experts (MoE) architectures. Experiments demonstrate its efficacy: by effectively selecting tokens, it substantially improves model performance across language, mathematics, reasoning, and coding tasks. The effectiveness is validated on both a small-scale dense model, Qwen3-1.7B, and a large-scale sparse MoE model, Qwen3-30B-A3B. Overall, our core contributions are:
@@ 2.1 Adaptive Token Selection
Token-level adaptive selection is most commonly applied in model quantization and compression (Yang et al., 2022b; a, Bondarenko et al., 2021), where Bayesian optimization methods are used to determine the appropriate compression ratios for tokens with varying levels of importance. This approach not only reduces computational redundancy but also mitigates the adverse effects of irrelevant tokens on the model’s attention mechanism. Beyond the realm of model compression, token selection has also been explored in the field of computer vision (Luo et al., 2025, Gadhikar et al., 2024, Hojjat et al., 2025). Motivated by the fact that visual representations often contain significant redundancy—due to irrelevant background information and high similarity between neighboring tokens—researchers have developed adaptive token selection strategies to address this issue. This strategy has been successfully applied to various vision tasks, including classification, detection, and retrieval. In general-purpose models, token selection is most notably employed in Mixture-of-Experts (MoE) architectures (Jiang et al., 2024, Yang et al., 2025, Liu et al., 2024), where a linear router dynamically assigns input tokens to specialized expert modules. Building on these insights, we propose the Dynamic Nested Depth (DND) approach, which adaptively selects tokens that are critical to represent and dynamically extends the model’s depth for their processing.
@@ 2.2 Dynamic Reasoning Depth
Dynamic adjustment mechanisms for inference paths can generally be classified into two primary approaches. The first line of work focuses on reducing computational redundancy, with notable techniques such as early exit (Zhou et al., 2025, Leviathan et al., 2023) and MOD (Raposo et al., 2024), which dynamically reduce the depth of computation layers to lower overall redundancy. The second line of research investigates the test-time scaling law, which shows that repeatedly processing tokens within a single layer can enhance final inference accuracy. A sophisticated variant of this approach is the Latent Strategy (Hao et al., 2024, Saunshi et al., 2025), where reasoning is carried out within hidden states—either by completing all steps before producing an answer or by leveraging recurrent inference to iteratively refine them.
The most closely related studies to our work are ITT (Chen et al., 2025) and MOR (Bae et al., 2025), both of which dynamically select a subset of tokens for additional computation and yield certain performance improvements. While sharing MOR (Bae et al., 2025)’s goal of improving performance via increased computational depth, our work differs in applicable model size, training phase, architecture, and routing control. MOR is limited to models with about 1B parameters and applied during pre-training, whereas our post-training approach scales from a 1.7B dense model to a 30B MoE model. Architecturally, MOR uses from-scratch training to treat dynamic depth output as the final state; in contrast, we employ a normalized fusion strategy to combine dynamic nested output with the original output state, thereby preserving previously learned knowledge. Furthermore, we identify and solve a key limitation in token selection control. Methods like MOR rely on z-loss (Zoph et al., 2022), which is inherently a load-balancing tool rather than a precise ratio controller. Therefore, we introduce a superior ratio control method. We enables precise and stable control over any predefined selection ratio target by simultaneously enhancing routing discriminability and adjusting the threshold through buffer error computation and EMA synchronization.
@@ 3.1 Architecture
In the model architecture section, we will introduce how tokens are selected (Sec. 3.1.1), how the new nested depth output is computed (Sec. 3.1.2), and how the vanilla output and dynamic nested output are fused to obtain the final output (Sec. 3.1.3). The whole architecture is shown in Fig. 2. Moreover, we apply the DND strategy only to the intermediate layers of the model, keeping the initial and final layers unchanged to preserve the reasoning patterns learned during pre-training (Ma et al., 2023, Xia et al., 2023). The layers where the DND is applied are denoted as from $\mathbf{L}_{s}$ to $\mathbf{L}_{e}$.
@@ 3.1.1 Routing Design
When considering the routing paradigm, as shown in Fig. 3 (a), routing the entire sequence with expert choice creates a mismatch with the next-token prediction paradigm of auto-regressive models. This is because the full sequence cannot be accessed during early decoding without risking information leakage (Raposo et al., 2024). To address this, we adopt token-choice selection, as illustrated in Fig. 3 (b), which dynamically decides whether each token should undergo further processing.
@@ 3.1.3 Fusion Design
Here, $\mathbf{x}^{i}$ refers to the merged hidden state of the $i$-th token. $\beta$ is a learnable parameter that acts as a balancing factor between the original and nested paths. Similar to (Raposo et al., 2024), this fusion is modulated by the token’s own routing score $p^{i}$, ensuring that tokens deemed more difficult (higher $p^{i}$) incorporate a larger portion of their recomputed representation. This design provides a smooth and adaptive integration, effectively stabilizing the learning process and allowing the model to control the influence of the additional computation dynamically.
@@ 3.2.2 Threshold Control Scheme
The preceding section introduced our method for enhancing the discriminability of router outputs. To further refine token selection, we propose a dynamic threshold control scheme, which adaptively regulates the threshold to achieve the desired proportion of selected tokens flexibly. Previous approaches using z-loss could only balance between selecting and not selecting tokens (Bae et al., 2025). Inspired by the balance loss proposed in DeepSeek-V3 (Liu et al., 2024), we design a method that computes the average selection ratio error in a buffer to adjust the threshold accordingly. Additionally, we leverage an EMA synchronization to assist in optimizing the threshold.
@@ 4.1 Evaluation Benchmark
We provide an extensive empirical evaluation of DND over a wide variety of benchmarks, demonstrating its effectiveness and robustness. The evaluation suite covers three primary domains: 1. General Knowledge & Alignment: MMLU (Hendrycks et al., 2020), CEval (Huang et al., 2023), CMMLU (Li et al., 2023), BBH (Srivastava et al., 2022), DROP (Dua et al., 2019), IFEval (Zhou et al., 2023), PIQA (Bisk et al., 2020). 2. Mathematics & STEM: Math (Hendrycks et al., 2021), GSM8k (Cobbe et al., 2021), MATH-500, AIME24, GPQA-Diamond (Rein et al., 2023),. 3. Coding & Agent: MBPP, MBPP+ (Austin et al., 2021), HumanEval+ (Chen et al., 2021), LCB-v5(LiveCodeBench-v5) (Jain et al., 2024), LCB-v6(LiveCodeBench-v6) (Jain et al., 2024), MultiPL-E (Cassano et al., 2022), BFCL v3 (Live).
@@ 4.2 Training Details
Our DND model undergoes supervised fine-tuning (SFT) using a comprehensive and diverse dataset. Our training data incorporates a significant volume of synthetic material built upon a high-quality seed set of 1-2 million instances curated from human annotations and open-source materials. The model’s weights are initialized from the Qwen3-30B-A3B Base model. Detailed hyperparameters and training settings are provided in Appendix Sec.B.
@@ 4.3 Main Results
As shown in Tab. 1, our DND strategy consistently improves the performance of the Qwen3-30B-A3B model, achieving an average gain of +0.87 across 17 benchmarks without any performance degradation. The impact of DND is most pronounced in Coding and Agent tasks, yielding notable gains of +2.05 on BFCL v3, +1.42 on LCB-v6, and +1.24 on LCB-v5. These results strongly support our hypothesis that DND effectively filters extraneous noise, allowing the model to focus its capacity on sparse, high-value tokens essential for complex reasoning, planning, and code generation. Importantly, the benefits are not limited to specialized domains: substantial improvements are also observed in General and Alignment tasks (+1.83 on C-Eval), alongside robust generalization on challenging Math and STEM benchmarks. Crucially, these substantial performance gains are realized with a negligible increase of only 0.03M parameters, highlighting DND as a highly parameter-efficient approach for unlocking significant capabilities in LLMs.
4.4 Ablation Study
DND Architecture. We first conducted an ablation study on the model’s architecture. The experimental results show that after applying our DND strategy, the SFT performance of Qwen3-1.7B improved significantly, with an average accuracy gain of 1.88 across tasks such as general, code, and math. Notably, there were substantial improvements on NLP tasks, such as BBH and GPQA-D, as well as on coding tasks. The results demonstrate the effectiveness of our DND architecture.
Training Strategies. We also conducted ablation experiments on the proposed training strategy. We found that when using only the DND framework with a simple z-loss-like method to control token selection, performance dropped noticeably, yielding an average improvement of only 1.01 points over Qwen3-1.7B’s SFT performance. This highlights the importance of our carefully designed training strategy. Furthermore, when testing a variant that only controls the router outputs or only dynamically adjusts the threshold, the performances similarly failed to reach the desired level.

### v2

@@ Abstract
We introduce Dynamic Nested Depth (DND), a novel method that improves performance for off-the-shelf LLMs by selecting critical tokens to reprocess in a nested depth manner. Specifically, at the end of the given transformer layer, DND identifies more critical tokens with a router and feeds them back for an extra round of processing, effectively “reviewing” difficult tokens while avoiding redundant computation for easier ones. The dynamic selection mechanism is tailored for precise control via two novel strategies: a router controlling loss to enhance token selection distinguishability, and a threshold control scheme to ensure selection stability. We demonstrate the effectiveness of DND by directly integrating it into pre-trained dense and MoE models during a post-training phase. On diverse benchmarks, DND boosts the performances of the dense Qwen3-1.7B, Llama3.2-1B, and Gemma3-1B by 1.88%, 2.61%, and 2.50% and the MoE Qwen3-30B-A3B by 0.87%, all with a minimal parameter and computing increase.
@@
Large Language Models (LLMs) have transformed artificial intelligence with their powerful abilities. The main strategy for improving them has been scaling, as empirical laws show that model performance predictably increases with more parameters, data, and computation (gpt, gemini, Yang2025Qwen3TR, DeepSeekAI2024DeepSeekV3TR). However, this scaling paradigm has significant drawbacks. The computational overhead for both training and inference grows exponentially with model size. This trend underscores a critical need for more efficient approaches to enhance model performance beyond simple brute-force scaling.
A key insight from (gloeckle2024better) is that prediction difficulty varies significantly across tokens; some are trivial to predict, while others demand deep computational processing. This disparity motivates token-level adaptive computation, where models can focus resources on the most critical inputs. A foundational version of this approach is token pruning, which has been shown to be effective across language understanding (mor), model compression (dta), and vision (vision3). By filtering out uninformative tokens, these methods reduce computational overhead and can even improve robustness by mitigating noise. This establishes a binary choice: a token is either discarded or processed normally. Furthermore, we propose the natural next step: instead of merely retaining challenging tokens for standard processing, we should allocate additional computation to them, ensuring these critical tokens are properly understood.
Our choice of how to allocate additional computation is inspired by latent strategies in test-time scaling (latent, saunshi2025reasoning). Unlike using explicit output expansion like COT, these methods recur computation in hidden states, scaling inference without extra text generation. They observed that reasoning tokens place uneven demands on computation: as illustrated in Fig. 1, most tokens serve fluency, while a few critical ones drive complex planning or logical transitions.
@@
To achieve dynamic selection and recalculation of critical tokens, we propose a novel architecture and training strategy. Specifically, as shown on the right of Fig. 1, we incorporate a linear layer as a router at the end of the transformer block with DND strategies. To achieve more robust routing and avoid potential information leakage during the inference of auto-regressive LLMs (MOD), we adopt a token-choice routing strategy. In this approach, each token is routed independently based on whether its output exceeds a predefined threshold. The selected tokens are reorganized into a new sequence and are re-fed to compute the dynamic nested output. Moreover, as DND is a post-training method, we carefully design a normalized fusion strategy that integrates the dynamic nested output with the original forward output to preserve global pre-training knowledge.
In terms of training strategies, since our routing method treats each token independently, it lacks the precision of the top-k ratio routing when selecting tokens for recurrent computation (MOD). To address this, we carefully design training strategies to control both the routers that determine token selection preferences and the thresholds that ultimately decide whether a token is selected. To enhance token selection distinguishability, we optimize the router’s output distribution, encouraging the outputs across tokens to be distinguishable through a router controlling loss. To stabilize the token selection ratio during training, we adopt a threshold control scheme, where the threshold is updated based on the error between the expected ratio and the actual ratio computed over a sample buffer. Furthermore, to ensure more synchronized control, we update the threshold by blending it with the average top-k routing values using an exponential moving average (EMA).
Our method is a post-training approach that can be directly integrated into existing dense and Mixture-of-Experts (MoE) architectures. Experiments demonstrate its efficacy: by effectively selecting tokens, it substantially improves model performance across language, mathematics, reasoning, and coding tasks. The effectiveness is validated on both three small-scale dense models, Qwen3-1.7B, Llama3.2-1B, Gemma3-1B and a large-scale sparse MoE model, Qwen3-30B-A3B. Overall, our core contributions are:
@@ 2.1 Adaptive Token Selection
Token-level adaptive selection is most commonly applied in model quantization and compression (dta, dqa, compress3), where Bayesian optimization methods are used to determine the appropriate compression ratios for tokens with varying levels of importance. This approach not only reduces computational redundancy but also mitigates the adverse effects of irrelevant tokens on the model’s attention mechanism. Beyond the realm of model compression, token selection has also been explored in the field of computer vision (vision1, vision2, vision3). Motivated by the fact that visual representations often contain significant redundancy—due to irrelevant background information and high similarity between neighboring tokens—researchers have developed adaptive token selection strategies to address this issue. This strategy has been successfully applied to various vision tasks, including classification, detection, and retrieval. In general-purpose models, token selection is most notably employed in Mixture-of-Experts (MoE) architectures (jiang2024mixtral, Yang2025Qwen3TR, DeepSeekAI2024DeepSeekV3TR), where a linear router dynamically assigns input tokens to specialized expert modules. Building on these insights, we propose the Dynamic Nested Depth (DND) approach, which adaptively selects tokens that are critical to represent and dynamically extends the model’s depth for their processing.
@@ 2.2 Dynamic Reasoning Depth
Dynamic adjustment mechanisms for inference paths can generally be classified into two primary approaches. The first line of work focuses on reducing computational redundancy, with notable techniques such as early exit (early1, early2) and MOD (MOD), which dynamically reduce the depth of computation layers to lower overall redundancy. The second line of research investigates the test-time scaling law, which shows that repeatedly processing tokens within a single layer can enhance final inference accuracy. A sophisticated variant of this approach is the Latent Strategy (latent, saunshi2025reasoning), where reasoning is carried out within hidden states—either by completing all steps before producing an answer or by leveraging recurrent inference to iteratively refine them.
The most closely related studies to our work are ITT (itt) and MOR (mor), both of which dynamically select a subset of tokens for additional computation and yield certain performance improvements. While sharing MOR (mor)’s goal of improving performance via dynamically increased computational depth, the two still differ fundamentally. MOR attempts to improve parameter efficiency during pretraining through a recurrent structure, which requires training a model from scratch on over 200B tokens. This is extremely costly and makes it difficult to apply the approach directly to existing open-source SOTA models. In contrast, DND focuses on unlocking the potential of existing state-of-the-art pretrained models and proposes a plug-and-play post-training method. Besides, our work differs from MOR in model scale, training phase, architecture, and routing control (detailed in Appendix. Sec . C). MOR is limited to 1B-parameter, whereas our DND successfully scales to a 30B MoE model. We also address a key limitation in token selection control. Unlike MOR, which relies on z-loss (zoph_2022_st) for approximate load balancing, we achieve precise, stable token selection. Our method jointly enhances routing discriminability and adjusts thresholds via EMA-synchronized buffer errors.
@@ 3.1 Architecture
In the model architecture section, we will introduce how tokens are selected (Sec. 3.1.1), how the new nested depth output is computed (Sec. 3.1.2), and how the vanilla output and dynamic nested output are fused to obtain the final output (Sec. 3.1.3). The whole architecture is shown in Fig. 2. Moreover, we apply the DND strategy only to the intermediate layers of the model, keeping the initial and final layers unchanged to preserve the reasoning patterns learned during pre-training (know1, know2). The layers where the DND is applied are denoted as from $\mathbf{L}_{s}$ to $\mathbf{L}_{e}$.
@@ 3.1.1 Routing Design
When considering the routing paradigm, as shown in Fig. 3 (a), routing the entire sequence with expert choice creates a mismatch with the next-token prediction paradigm of auto-regressive models. This is because the full sequence cannot be accessed during early decoding without risking information leakage (MOD). To address this, we adopt token-choice selection, as illustrated in Fig. 3 (b), which dynamically decides whether each token should undergo further processing.
@@ 3.1.3 Fusion Design
Here, $\mathbf{x}^{i}$ refers to the merged hidden state of the $i$-th token. $\beta$ is a learnable parameter that acts as a balancing factor between the original and nested paths. Similar to (MOD), this fusion is modulated by the token’s own routing score $p^{i}$, ensuring that tokens deemed more difficult (higher $p^{i}$) incorporate a larger portion of their recomputed representation. This design provides a smooth and adaptive integration, effectively stabilizing the learning process and allowing the model to control the influence of the additional computation dynamically.
@@ 3.2.2 Threshold Control Scheme
The preceding section introduced our method for enhancing the discriminability of router outputs. To further refine token selection, we propose a dynamic threshold control scheme, which adaptively regulates the threshold to achieve the desired proportion of selected tokens flexibly. Previous approaches using z-loss could only balance between selecting and not selecting tokens (mor). Inspired by the balance loss proposed in DeepSeek-V3 (DeepSeekAI2024DeepSeekV3TR), we design a method that computes the average selection ratio error in a buffer to adjust the threshold accordingly. Additionally, we leverage an EMA synchronization to assist in optimizing the threshold.
@@ 4.1 Evaluation Benchmark
We provide an extensive empirical evaluation of DND over a wide variety of benchmarks, demonstrating its effectiveness and robustness. The evaluation suite covers three primary domains: 1. General Knowledge & Alignment: MMLU (mmlu), CEval (huang_2023_ceval), CMMLU (li2023cmmlu), BBH (bbh), DROP (DROP), IFEval (ifeval), PIQA (bisk2020piqa). 2. Mathematics & STEM: Math (math), GSM8k (cobbe2021gsm8k), MATH-500, AIME24, GPQA-Diamond (Rein2023GPQAAG),. 3. Coding & Agent: MBPP, MBPP+ (mbpp), HumanEval+ (humaneval), LCB-v5(LiveCodeBench-v5) (Jain2024LiveCodeBenchHA), LCB-v6(LiveCodeBench-v6) (Jain2024LiveCodeBenchHA), MultiPL-E (Cassano2022MultiPLEAS), BFCL v3 (Live).
@@ 4.2 Training Details
Our DND model undergoes standard full-scale supervised fine-tuning (SFT) using a comprehensive and diverse dataset, with all parameters set as trainable and the same learning rate applied. Our training data incorporates a significant volume of synthetic material built upon a high-quality seed set of 1-2 million instances curated from human annotations and open-source materials. The model’s weights are initialized from the Qwen3-1.7B Base, Llama-3.2-1B, Gemma3-1B-pt, and Qwen3-30B-A3B Base. Detailed hyperparameters and training settings are provided in Appendix Sec.B.
@@ 4.3 Main Results
Table 1: SFT Performance Comparison of Different Small-Scale Dense LLMs. Performing full-scale SFT with the DND strategy on the three widely used base models yields additional average improvements of 1.88, 2.61, and 2.50 points over full-scale SFT alone.
General Knowledge & Alignment Math & STEM Coding & Agent
Average BBH PIQA C-Eval MMLU IFEval GPQA GSM8K MBPP Human BFCL MultiPLE
Qwen3-1.7B 59.53 40.82 75.38 60.00 64.11 65.47 28.54 79.38 68.38 61.59 58.73 52.45
+ ITT 59.58 41.23 75.99 59.51 64.92 65.08 27.85 80.29 68.67 60.78 58.62 52.45
+ DND 61.41 45.84 76.25 60.38 64.45 66.87 34.34 80.15 71.90 62.71 59.80 52.80
$\Delta$ (+-) +1.88 +5.02 +0.87 +0.38 +0.34 +1.40 +5.80 +0.77 +3.52 +1.12 +1.07 +0.35
Llama3.2-1B 45.37 25.73 65.48 47.82 53.28 52.45 10.73 63.23 49.54 50.42 44.89 35.47
+ DND 47.98 29.43 66.51 49.21 55.63 55.68 14.59 66.57 52.91 52.16 47.52 37.56
$\Delta$ (+-) +2.61 +3.70 +1.03 +1.39 +2.35 +3.23 +3.86 +3.34 +3.37 +1.74 +2.63 +2.09
Gemma3-1B 47.08 25.93 70.27 50.14 55.98 54.52 16.49 65.53 49.29 52.62 40.69 36.43
+ DND 49.58 30.62 71.33 51.00 58.04 56.91 21.79 68.68 52.96 53.76 42.88 37.41
$\Delta$ (+-) +2.50 +4.69 +1.06 +0.86 +2.06 +2.39 +5.30 +3.15 +3.67 +1.14 +2.19 +0.98
Base Evaluation. As shown in Tab. 1, our method achieves obvious improvements across the three widely used base models. Especially on datasets that require complex reasoning, such as BBH and GPQA, the performance boost is particularly notable, with all three models showing an additional performance improvement of around 5%. Additionally, we found that when the SFT is conducted with ITT (itt) under the same computation cost, the performance improvement is not as pronounced. The limited performance gains stem from the use of Top-P–based token selection for auto-regressive LLM, which introduces a mismatch between training and inference, and may also lead to potential information leakage according to (MOD).
@@
Scaling Evaluation. As shown in Tab. 2, our DND strategy consistently improves the performance of the Qwen3-30B-A3B model, achieving an average gain of +0.87 across 17 benchmarks without any performance degradation. The impact of DND is most pronounced in Coding and Agent tasks, yielding notable gains of +2.05 on BFCL v3, +1.42 on LCB-v6, and +1.24 on LCB-v5. These results strongly support our hypothesis that DND effectively filters extraneous noise, allowing the model to focus its capacity on sparse, high-value tokens essential for complex reasoning, planning, and code generation. Importantly, the benefits are not limited to specialized domains: substantial improvements are also observed in General and Alignment tasks (+1.83 on C-Eval), alongside robust generalization on challenging Math and STEM benchmarks. Crucially, these substantial performance gains are realized with a negligible increase of only 0.03M parameters, highlighting DND as a highly parameter-efficient approach for unlocking significant capabilities in LLMs. FLOPs and Throughput Evaluation. As demonstrated in Appendix Sec. A, reviewing 20% of the tokens adds only about 6% extra FLOPs when applying the DND strategy to the Qwen3-30B-A3B model. To further assess the inference speed of our model in practical settings relative to the baseline, following SGLang, as shown in the Tab. 3, we measured the throughput of both Qwen3-30B-A3B and Qwen3-30B-A3B+DND models under four standard sequence lengths using a single H100 GPU with a single batch. The measurement results in the table show that, while achieving performance improvements, our model consistently reaches 91.6-93.1% of the speed of the vanilla model under different circumstances.
Table 3: Comparison of Speed across Different Input and Decode Lengths. The speeds are measured using BF16 Quantization of LLM and accelerated with vLLM kernels.
   Input Length    Decode Length    Speed (tokens/s)    Relative Speed
   Qwen3-30B-A3B    + DND    (%)
   1024    2048    148.68    136.19    91.6
   6144    2048    208.60    193.51    92.8
   1024    6144    76.16    70.52    92.6
   6144    6144    100.89    93.93    93.1
@@
(a) Entropy of Selected Tokens. The selection frequency of a token correlates positively with its original logit entropy in the vanilla pass.
(b) Logits Variation after DND. For rarely selected tokens, logit entropy fluctuates evenly, but it decreases as selection frequency increases.
4.4 Ablation Study
Training Strategies. We conducted ablation experiments on the proposed training strategy. We found that when using only the DND framework with a simple z-loss-like method to control token selection, performance dropped noticeably, yielding an average improvement of only 1.01 points over Qwen3-1.7B’s SFT performance. This highlights the importance of our carefully designed training strategy. Moreover, the router and threshold control function as complementary components for token selection control. While each method individually provides marginal gains, their combination leads to a clear improvement of approximately one percentage point in average accuracy.